OpenStudy (bluebeta):

PLEASE HELP!! For each complex number below, compute the modulus and argument. 1. -2i 2.-5 3.5-5i 4. 7sqrt3+7i 5.-3+4i

OpenStudy (bluebeta):

@mathmale @surjithayer

OpenStudy (anonymous):

\[|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

although you don't need that for the first one, since the absolute value of \(-2i\) is obvious without computation

OpenStudy (bluebeta):

@satellite73 what do you mean? I know I probably sound stupid but I'm really confused.

OpenStudy (anonymous):

\(-2i\) is here 

OpenStudy (anonymous):

how far is that away from the origin?

OpenStudy (anonymous):

if you are really confused, i will tell you, but it should be pretty clear how many using that is away from the origin

OpenStudy (anonymous):

"units" i meant

OpenStudy (anonymous):

if you know, let me know if not, let me know an i will tell you

OpenStudy (bluebeta):

its 2 units down

OpenStudy (anonymous):

yes, but of course absolute value is positive, don't worry about the direction so the "modulus" is 2

OpenStudy (anonymous):

what's next?

OpenStudy (bluebeta):

so r=2?

OpenStudy (anonymous):

yes, if you are calling is \(r\) then \(r=2\)

OpenStudy (bluebeta):

ok now the argument?

OpenStudy (anonymous):

that is the angle it makes with the positive real line (like the x axis) before we begin are you working in degrees, like \(45\) and \(90\) or in radians, like \(\frac{\pi}{4}\) and \(\frac{\pi}{2}\)?